OpenStudy (anonymous):

A cone has a diameter 8 iin. and height 14in. A rectangular prism is 6 in. by 4in. by 10 in. A square pyramid has base edge 8in and height 12 in. What are the volumes of the three figures in order from least to greatest. A. cone, prism,pyramid B. prism, cone, pyramind C. pyramid, cone, prism D. prism, pyramid, cone 3. what is the diameter of a sphere whose surphace area is 100 pi m ^2 ? A.5m B.10m C.5 pi m D.25 pi m

OpenStudy (anonymous):

Do you know the formulas for the volume of a cone, of a prism, and of a pyramid?

OpenStudy (anonymous):

no .

OpenStudy (anonymous):

Mkay, then lets solve each one individually. Here's the reference sheet for future reference. http://fcat.fldoe.org/eoc/pdf/algebra1.pdf

OpenStudy (anonymous):

So lets do the cone. V = 1/3Bh Capital B = area of the base.

OpenStudy (anonymous):

okay .thanks

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

formulas are: Cone- V=1/3(pi)*r^2*h Rec. Prism- V=l*w*h Sq. Pyramid- V=1/3*b^2*h

OpenStudy (anonymous):

So find B first. Area of a circle is the radius^2 times pi. B = 4^2 * 3.14 B = 16 * 3.14 B = 50.24 And our height is 14. V = 1/3Bh V = 1/3(50.24)(14)

OpenStudy (anonymous):

waiitt... Jazy's got it!(:

OpenStudy (anonymous):

Or \[V = \frac{ 50.24 \times 14}{ 3 }\]Can you do that with a calculator?

OpenStudy (anonymous):

yeah . 234.4

OpenStudy (anonymous):

Alright. So the Volume for the cone is 234.4. Now the rectangular prism. 6 in 4 in 10 in V = bwh V = (6)(4)(10) V = ?

OpenStudy (anonymous):

240

OpenStudy (anonymous):

Perfect. Now the Square pyramid: base = 8 height = 12 V = 1/3Bh V = 1/3B(12) Find B: Its a square so to find the area multiply 8*8 = 64. V = 1/3(64)(12) \[V = \frac{ 64 \times 12 }{ 3 }\]\[\large V =~ ?\]

OpenStudy (anonymous):

256

OpenStudy (anonymous):

cone, prism, pyramid

OpenStudy (anonymous):

d(^.^)b

OpenStudy (anonymous):

what ?

OpenStudy (anonymous):

Its a thumbs up. lol

OpenStudy (anonymous):

oh okay lol . Next question please .

sam (.sam.):

number 3?

OpenStudy (anonymous):

yes .

OpenStudy (aravindg):

surface area of sphere=\(4 \pi r^2\) r=d/2

OpenStudy (aravindg):

you are given the syrface area, Solve for d .

sam (.sam.):

Surface area of sphere = \(4\pi r^2\) \[A=4 \pi r^2 \\ \\ 100\pi=4\pi r^2 \\ \\ r^2=\frac{100}{4} \\ \\ r=\frac{10}{2} \\ \\ r=5 \\ \\ \] Diameter is twice of radius

OpenStudy (anonymous):

thank you

sam (.sam.):